OpenStudy (anonymous):

Jon's new mini computer monitor has a 20 cm by 24 cm LCD screen with a plastic trim border. The total area of the monitor and border is 672 cm2. How wide is the trim around the monitor?

OpenStudy (anonymous):

(20+x)(24+x)=672 x is the width of the border

OpenStudy (anonymous):

I got x^2+44x+480=672. What do I do now?

OpenStudy (anonymous):

solve x^2+44x-192=0

OpenStudy (anonymous):

How do I solve that?

OpenStudy (anonymous):

you can factor what factors of -192 add up to +44, (one is negative, one is positive)

OpenStudy (anonymous):

+4 & -48.

OpenStudy (anonymous):

close, +48 and -4

OpenStudy (anonymous):

so your factored equation is (x+48)(x-4)=0

OpenStudy (anonymous):

x=-48, 4 because length can't be negative, you discard the -48 the width of the border is 4 cm

OpenStudy (anonymous):

you mean, you discard -4 and have a width of 48 cm